# Intensity-Efficient Allocations

### v5

@@ 1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
For the “only if” part, consider an arbitrary Pareto-efficient allocation $x$. Suppose to the contrary that (i) is false. That is, $x_{j}\neq b_{j}^{1}$ for some $v_{j}\in V^{1}$. Let $b_{j}^{1}=x_{l}$ for the relevant distinct agent $v_{l}\in V$ (such an agent exists because $|V|=|A|$). Either $v_{l}\in V^{1}$ or $v_{l}\in V^{i>1}$. In the former case, swapping $x_{j}$ and $x_{l}$ between the two agents Pareto-improves upon $x$, contradicting its postulated Pareto optimality. In the latter case, from the assumption that $V^{i>1}$ is top-preference-coincident with respect to $A^{i>1}$, and that $S$ is homeo-monotonic, we have $a^{i}_{t}\succ_{l}x_{l}\equiv b^{1}_{j}$ for every $a^{i}_{t}\in A^{i>1}$. Furthermore, in this case (again because $|V|=|A|$) there must also exist another agent $k\in V^{p}$, $p\neq i$, such that $x_{k}\in A^{i>1}$. There are two possibilities: $p=1$ or $p>1$. If $p=1$, then the above contradiction is obtained now for $x_{l}$ and $x_{k}$. If $p>1$, then, by homeo-monotonicity of $S$ and the implied structure of $(V^{l>1},A^{l>1})$, $(V^{p>1},A^{p>1})$, we have $s_{l}(x_{k},x_{l})>0$ and $s_{k}(x_{l},x_{k})>0$, which also contradicts the postulated Pareto-efficiency of $x$. The argument establishing necessity of (ii) is analogous and omitted. ∎
@@
Next, consider a top-preference-coincident set $V^{i>1}$ with respect to $A^{i>1}$. By effective homeo-monotonicity, either such a non-empty set exists or the envy-free set $V^{1}$ coincides with $V$. In the latter case, the allocation $x^{*}\in\mathcal{A}$ defined by $x^{*}_{i}=b^{1}_{i}$ is obviously the unique Pareto-efficient allocation at $S$ and therefore, trivially, intensity-efficient. So, in the sequel we assume that $V^{i>1}\neq\emptyset$.
@@
Denote by $a_{1}^{i}\succ_{i}a_{2}^{i}\succ_{i}\cdots\succ_{i}a_{|A^{i}|}^{i}$ the induced preference order on $A^{i>1}$ which, by the definition of $(V^{i>1},A^{i>1})$, is common across all agents in $V^{i>1}$. Define the binary relation $\trianglerighteq^{i}$ on $V^{i>1}$ by
@@
By the postulated monotonicity of $(V^{i>1},A^{i>1})$, $\trianglerighteq^{i}$ is a weak order on $V^{i>1}$. Let $\trianglerighteq_{i}^{*}$ be some extension of $\trianglerighteq^{i}$ into a linear order over $V^{i>1}$, and denote by $\trianglerighteq^{*i}(j)$ the $\trianglerighteq^{*i}$-rank of agent $v_{j}\in V^{i>1}$. By construction, $\trianglerighteq^{*i}$ coincides with $\trianglerighteq^{i}$ except where it breaks any equivalence ties that the latter relation may feature.
@@
$x_{j}^{*1}:=b_{j}^{1},\;j=1,\ldots,|V^{1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{1}$
$x_{j}^{*i>1}:=a^{i}_{l},\;j=1,\ldots,|V^{i>1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{i>1}\text{ and }\trianglerighteq^{*i}(j)=l$
That is, each agent in the set $V^{1}$ (which could be empty) is assigned their most preferred item, while the $\trianglerighteq^{*i}$-$l$th-ranked agent in $V^{i>1}$ is assigned their $l$th most preferred item. Recalling that $V^{1}$ is envy-free and each $V^{i>1}$ is top-preference-coincident, it follows from Lemma 3 that $x^{*}$ is Pareto-efficient. We will prove that it is also intensity-efficient.
To this end, suppose to the contrary that some Pareto-efficient allocation $y\in\mathcal{A}$ intensity-dominates $x^{*}$. Then, there are agents $v_{i},v_{j}\in V$ such that $(x_{i}^{*},x_{j}^{*})=(y_{j},y_{i})$ and $s_{j}(y_{j},y_{i})\equiv s_{j}(x_{i}^{*},x_{j}^{*})>s_{i}(x^{*}_{i},x_{j}^{*})>0$. By construction of $x^{*}$, $v_{i},v_{j}$ cannot belong to the same $V^{l\geq 1}$. For $l=1$ this is obvious because $x^{*i}=b^{1}_{i}$ and $x^{*j}=b^{1}_{j}$, contradicting the above postulate. For $l>1$, moreover, one has $x^{*}_{i},x_{j}^{*}\in A^{l>1}$ while items in $A^{l>1}$ are assigned to agents in $V^{l>1}$ according to the linear order $\trianglerighteq^{*l}$, which, by homeo-monotonicity of $S$, ensures that $s_{i}(x_{i}^{*},x_{j}^{*})\geq s_{j}(x_{i}^{*},x_{j}^{*})$, thereby contradicting that postulate. The remaining possibility is $v_{i}\in V^{l>1}$ and $v_{j}\in V^{k>1}$ for some $k\neq l$. By construction of $x^{*}$, however, we have $x_{i}^{*}\in A^{l>1}$ and $x_{j}^{*}\in A^{k>1}$ which, because $V^{l>1}$, $V^{k>1}$ are top-preference-coincident with respect to $A^{l>1}$, $A^{k>1}$ and $A^{l>1}\cap A^{k>1}=\emptyset$, implies $s_{i}(x_{i}^{*},x_{j}^{*})\equiv s_{i}(y_{j},y_{i})>0$ and $s_{j}(x_{i}^{*},x_{j}^{*})\equiv s_{j}(y_{j},y_{i})<0$. Thus, the postulate that $y$ intensity-dominates $x^{*}$ is contradicted in all situations. ∎
Example 2 (continued).
The way in which the intensity-efficient allocation $x^{*}$ is constructed in the proof of Theorem 2 readily suggests the following stronger implication for a class of special cases:
Corollary 2.
An intensity-dominant—up to equivalence—allocation exists in any effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where every top-preference-coincident set has two agents.
Importantly, although effective homeo-monotonicity is a sufficient but not necessary condition for the existence of intensity-efficient allocations, we show next that the restriction it imposes within monotonic top-preference-coincident sets—if the profile induces such sets—is also necessary whenever these sets are strictly monotonic.
Theorem 3.
Proof.
@@
The only difference between allocations $x$ and $x^{\prime}$ lies in how they permute items in $A^{\prime}$ to agents in $V^{\prime}$. Denote their restrictions on these subsets by $\widehat{x}$ and $\widehat{x}^{\prime}$, respectively. By Lemma 3 and the postulated intensity-efficiency (hence Pareto-efficiency) of $x$ it follows that $x^{\prime}$ is Pareto-efficient too. Since $x$ is intensity-efficient, it is not intensity-dominated by $x^{\prime}$. Hence, there is no pair $(a^{\prime},a^{\prime\prime})\in A^{\prime}\times A^{\prime}$ such that $(a^{\prime},a^{\prime\prime})=(\widehat{x}_{i},\widehat{x}_{j})=(\widehat{x}^{\prime}_{j},\widehat{x}^{\prime}_{i})$ for some $v^{\prime}_{i},v^{\prime}_{j}\in V^{\prime}$, making $x$ and $x^{\prime}$ incomparable by intensity-dominance. The argument in the proof of Theorem 2 following Lemma 3 establishes that there is no other Pareto-efficient allocation that intensity-dominates $x^{\prime}$ either. Thus, $x^{\prime}$ is intensity-efficient. ∎
@@
$\begin{array}[]{lllllllll}s^{\prime}_{1}(a^{\prime}_{i},a^{\prime}_{i+1})&>&s^{\prime}_{2}(a^{\prime}_{i},a^{\prime}_{i+1})&>&\cdots&>&s^{\prime}_{|V^{\prime}|}(a^{\prime}_{i},a^{\prime}_{i+1})&&\text{for all }i\leq|A^{\prime}|-1\end{array}$
Now recall that the allocation $x^{\prime}$ constructed above is intensity-efficient. This assigns $x^{\prime}_{m}=a^{\prime}_{m}\in A^{\prime}$ to each $v^{\prime}_{m}\in V^{\prime}$. Since $(V^{\prime},A^{\prime})$ is strictly monotonic but not homeo-monotonic, there are $v^{\prime}_{j},v^{\prime}_{l}\in V^{\prime}$ such that $v^{\prime}_{j}\trianglerighteq^{\prime}v^{\prime}_{l}$, jl>1j-l>1 and $s^{\prime}_{l}(a^{\prime}_{j},a^{\prime}_{l})>s^{\prime}_{j}(a^{\prime}_{j},a^{\prime}_{l})$. Since $(x^{\prime}_{j},x^{\prime}_{l})=(a^{\prime}_{j},a^{\prime}_{l})$, this implies that the allocation $x^{\prime\prime}$ which is identical to $x^{\prime}$ except that $(x^{\prime\prime}_{j},x^{\prime\prime}_{l})=(a^{\prime}_{l},a^{\prime}_{j})$ intensity-dominates $x^{\prime}$, contradicting the postulated intensity-efficiency of the latter allocation. ∎
Example 3.
For $n=4$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as follows:
$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$
1 2 3 4
6 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{4},a_{1})$
5 $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{1},a_{3})$
4 $(a_{1},a_{3})$ $(a_{1},a_{3})$ $(a_{2},a_{4})$
3 $(a_{1},a_{2})$ $(a_{1},a_{2})$ $(a_{3},a_{4})$
2 $(a_{2},a_{3})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
1 $(a_{3},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$
The Pareto-efficient allocations here are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$. Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$ and $z$ (these are dominated by $p$ and $y$, respectively). Yet $V^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$. $\blacklozenge$
Example 4.
For $n=5$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4},v_{5}\},\{a_{4},a_{5}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\}$, $\{a_{1},a_{2},a_{3}\})$, as follows:
$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{1}$\hskip 20.0pt}}$
1 2 3    4    5
10 $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{4},\cdot)$ $(a_{5},\cdot)$
9 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $\vdots$ $\vdots$
8 $(a_{2},a_{5})$ $(a_{2},a_{5})$ $(a_{1},a_{3})$
7 $(a_{2},a_{4})$ $(a_{1},a_{3})$ $(a_{2},a_{5})$
6 $(a_{1},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$
5 $(a_{3},a_{5})$ $(a_{3},a_{5})$ $(a_{3},a_{5})$
4 $(a_{1},a_{2})$ $(a_{3},a_{4})$ $(a_{3},a_{4})$
3 $(a_{2},a_{3})$ $(a_{1},a_{2})$ $(a_{4},a_{5})$
2 $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
1 $(a_{4},a_{5})$ $(a_{4},a_{5})$ $(a_{2},a_{3})$
The Pareto-efficient allocations here are
Yet the intensity-dominance relation $D$ is cyclic over this set:
@@ 1 $(a_{4},a_{5})$ $(a_{4},a_{5})$ $(a_{2},a_{3})$
$\blacklozenge$
Together, Theorems 2-3 and Examples 2-4 highlight the importance of the intuitive structure imposed by homeo-monotonicity for the general existence of intensity-efficient allocations, both as a sufficient condition and, in a weaker sense, as a necessary one too. As the counterexample for $n=4$ in the proof of Theorem 1 shows, moreover, a profile may fail to have an intensity-efficient allocation even when homeo-monotonicity as defined above is satisfied, yet a subtle variation of it is not. More specifically, that profile features: (i) $n=4$; (ii) $a\succ_{i}b\succ_{i}c\succ_{i}d$ for $i\leq 4$ (hence a top-preference-coincident set $V$ with respect to $A$); (iii) $s_{1}=s_{4}$, $s_{2}=s_{3}$. Although the four agents’ intensities over their common preferences do not define a monotonic top-preference-coincident set, one observes that emergent here is a variant of this notion whereby each set in the partition $\{\{v_{1},v_{2}\},\{v_{3},v_{4}\}\}$ of $V$ is monotonically top-preference-coincident with respect to $\{a,b,c\}\subset A$: $s_{1}(a,b)>s_{2}(a,b)$, $s_{1}(b,c)=s_{2}(b,c)$ and $s_{4}(a,b)>s_{3}(a,b)$, $s_{4}(b,c)=s_{3}(b,c)$. However, both these pairs of agents violate what homeo-monotonicity would require here: $s_{1}(a,c)<s_{2}(a,c)$ and $s_{4}(a,c)<s_{3}(a,c)$. Thus, while the problem of fully characterizing the conditions under which strict profiles have intensity-efficient allocations remains open, the preceding analysis hopefully provides non-trivial and useful guidance in that direction.
Figure 1: Simulations-based histograms showing the proportions of Pareto-efficient allocations that are also intensity-efficient when the number of agents and items is small.
Additional motivation for further study is also contributed by the analysis of uniform-randomly drawn intensity profiles when $n=3,4,5$. Figure 1 summarizes this output and displays how the distribution of the per-profile proportion of Pareto-efficient allocations that are also intensity-efficient varies with $n$. Interestingly, while the average number of Pareto-efficient allocations increases rapidly from 2.8 to 5.9 to 14.1 when $n=3,4,5$, respectively, the corresponding averages for intensity-efficient allocations are 1.6, 1.9 and 2.4. The refinement gains afforded by intensity-efficiency are therefore significant and appear to be increasing in $n$, with the mean proportion of intensity-efficient out of all Pareto-efficient allocations falling from 0.63 ($n=3$) to 0.2 ($n=5$). Finally, non-existence of intensity-efficient allocations is a rare occurrence in these simulations, with only 1 and 15 out of the respective 100,000 randomly drawn profiles being associated with a cyclic intensity-dominance relation over Pareto-efficient allocations when $n=4$ and $n=5$, respectively.
We end this section by remarking that the paper’s new analytical environment necessitated novel arguments in the proofs of this section’s main results. More specifically, viewed as an abstract mathematical problem, existence of intensity-efficient allocations amounts to identifying sufficient conditions for acyclicity of the intensity-dominance binary relation defined on the set of permutations over alternatives. The mathematical literature on the combinatorics of permutations has studied partial orders over such sets, which may even have the richer structure of a lattice that additionally ensures the existence of both a greatest (hence maximal) and a smallest element. These results, however, are not applicable in our environment. The reason is that those partial orders, namely the strong and weak Bruhat orders (Stanley, 2012, pp. 399-400), relate two permutations if and only if there is a single transposition (“flip”) between elements in the two permutations, and this transposition corresponds to an inversion (Stanley, 2012, p. 30) relative to some postulated fixed linear order over its elements. The weak Bruhat order (see Ceballos and Pons (2024) for a recent generalization) gives rise to a lattice—whose graph is known as the permut(a)(o)hedron—but under the additional assumption that the said single transposition occurs between adjacent elements of the two relevant permutations. Our problem on the other hand does not assume a fixed linear order over items and is defined by intensity-dominance comparisons that generally feature possibly multiple transpositions of non-adjacent elements, which are defined by dominance between specific entries in the two relevant columns of an integer matrix, defined in turn by the corresponding agents’ intensity orderings.
5 Eliciting Ordinal Intensities and Finding Intensity-Dominating Allocations
We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems. In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency in the intensity-dominating direction. Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational. To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
Proposition 3.
Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in exactly $\frac{1}{2}(n-1)(n-2)$ steps, hence in $O(n^{2})$ time.

### v6

@@ 6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
For the “only if” part, consider an arbitrary Pareto-efficient allocation $x$. Suppose to the contrary that (i) is false. That is, $x_{j}\neq b_{j}^{1}$ for some $v_{j}\in N^{1}$. Let $b_{j}^{1}=x_{l}$ for the relevant distinct agent $v_{l}\in N$ (such an agent exists because $|N|=|A|$). Either $v_{l}\in N^{1}$ or $v_{l}\in N^{i>1}$. In the former case, swapping $x_{j}$ and $x_{l}$ between the two agents Pareto-improves upon $x$, contradicting its postulated Pareto optimality. In the latter case, from the assumption that $N^{i>1}$ is top-preference-coincident with respect to $A^{i>1}$, and that $S$ is homeo-monotonic, we have $a^{i}_{t}\succ_{l}x_{l}\equiv b^{1}_{j}$ for every $a^{i}_{t}\in A^{i>1}$. Furthermore, in this case (again because $|V|=|A|$) there must also exist another agent $k\in V^{p}$, $p\neq i$, such that $x_{k}\in A^{i>1}$. There are two possibilities: $p=1$ or $p>1$. If $p=1$, then the above contradiction is obtained now for $x_{l}$ and $x_{k}$. If $p>1$, then, by homeo-monotonicity of $S$ and the implied structure of $(N^{l>1},A^{l>1})$, $(N^{p>1},A^{p>1})$, we have $s_{l}(x_{k},x_{l})>0$ and $s_{k}(x_{l},x_{k})>0$, which also contradicts the postulated Pareto-efficiency of $x$. The argument establishing necessity of (ii) is analogous and omitted. ∎
@@
Next, consider a top-preference-coincident set $N^{i>1}$ with respect to $A^{i>1}$. By effective homeo-monotonicity, either such a non-empty set exists or the envy-free set $N^{1}$ coincides with $N$. In the latter case, the allocation $x^{*}\in\mathcal{A}$ defined by $x^{*}_{i}=b^{1}_{i}$ is obviously the unique Pareto-efficient allocation at $S$ and therefore, trivially, intensity-efficient. So, in the sequel we assume that $N^{i>1}\neq\emptyset$.
@@
Denote by $a_{1}^{i}\succ_{i}a_{2}^{i}\succ_{i}\cdots\succ_{i}a_{|A^{i}|}^{i}$ the induced preference order on $A^{i>1}$ which, by the definition of $(N^{i>1},A^{i>1})$, is common across all agents in $N^{i>1}$. Define the binary relation $\trianglerighteq^{i}$ on $N^{i>1}$ by
@@
By the postulated monotonicity of $(N^{i>1},A^{i>1})$, $\trianglerighteq^{i}$ is a weak order on $N^{i>1}$. Let $\trianglerighteq_{i}^{*}$ be some extension of $\trianglerighteq^{i}$ into a linear order over $N^{i>1}$, and denote by $\trianglerighteq^{*i}(j)$ the $\trianglerighteq^{*i}$-rank of agent $v_{j}\in N^{i>1}$. By construction, $\trianglerighteq^{*i}$ coincides with $\trianglerighteq^{i}$ except where it breaks any equivalence ties that the latter relation may feature.
@@
$x_{j}^{*1}:=b_{j}^{1},\;j=1,\ldots,|N^{1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in N^{1}$
$x_{j}^{*i>1}:=a^{i}_{l},\;j=1,\ldots,|N^{i>1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in N^{i>1}\text{ and }\trianglerighteq^{*i}(j)=l$
That is, each agent in the set $N^{1}$ (which could be empty) is assigned their most preferred item, while the $\trianglerighteq^{*i}$-$l$th-ranked agent in $N^{i>1}$ is assigned their $l$th most preferred item. Recalling that $N^{1}$ is envy-free and each $N^{i>1}$ is top-preference-coincident, it follows from Lemma 3 that $x^{*}$ is Pareto-efficient. We will prove that it is also intensity-efficient.
To this end, suppose to the contrary that some Pareto-efficient allocation $y\in\mathcal{A}$ intensity-dominates $x^{*}$. Then, there are agents $v_{i},v_{j}\in N$ such that $(x_{i}^{*},x_{j}^{*})=(y_{j},y_{i})$ and $s_{j}(y_{j},y_{i})\equiv s_{j}(x_{i}^{*},x_{j}^{*})>s_{i}(x^{*}_{i},x_{j}^{*})>0$. By construction of $x^{*}$, $v_{i},v_{j}$ cannot belong to the same $N^{l\geq 1}$. For $l=1$ this is obvious because $x^{*i}=b^{1}_{i}$ and $x^{*j}=b^{1}_{j}$, contradicting the above postulate. For $l>1$, moreover, one has $x^{*}_{i},x_{j}^{*}\in A^{l>1}$ while items in $A^{l>1}$ are assigned to agents in $N^{l>1}$ according to the linear order $\trianglerighteq^{*l}$, which, by homeo-monotonicity of $S$, ensures that $s_{i}(x_{i}^{*},x_{j}^{*})\geq s_{j}(x_{i}^{*},x_{j}^{*})$, thereby contradicting that postulate. The remaining possibility is $v_{i}\in N^{l>1}$ and $v_{j}\in N^{k>1}$ for some $k\neq l$. By construction of $x^{*}$, however, we have $x_{i}^{*}\in A^{l>1}$ and $x_{j}^{*}\in A^{k>1}$ which, because $N^{l>1}$, $N^{k>1}$ are top-preference-coincident with respect to $A^{l>1}$, $A^{k>1}$ and $A^{l>1}\cap A^{k>1}=\emptyset$, implies $s_{i}(x_{i}^{*},x_{j}^{*})\equiv s_{i}(y_{j},y_{i})>0$ and $s_{j}(x_{i}^{*},x_{j}^{*})\equiv s_{j}(y_{j},y_{i})<0$. Thus, the postulate that $y$ intensity-dominates $x^{*}$ is contradicted in all situations. $\square$
@@
The only difference between allocations $x$ and $x^{\prime}$ lies in how they permute items in $A^{\prime}$ to agents in $N^{\prime}$. Denote their restrictions on these subsets by $\widehat{x}$ and $\widehat{x}^{\prime}$, respectively. By Lemma 3 and the postulated intensity-efficiency (hence Pareto-efficiency) of $x$ it follows that $x^{\prime}$ is Pareto-efficient too. Since $x$ is intensity-efficient, it is not intensity-dominated by $x^{\prime}$. Hence, there is no pair $(a^{\prime},a^{\prime\prime})\in A^{\prime}\times A^{\prime}$ such that $(a^{\prime},a^{\prime\prime})=(\widehat{x}_{i},\widehat{x}_{j})=(\widehat{x}^{\prime}_{j},\widehat{x}^{\prime}_{i})$ for some $N^{\prime}_{i},N^{\prime}_{j}\in N^{\prime}$, making $x$ and $x^{\prime}$ incomparable by intensity-dominance. The argument in the proof of Theorem 2 following Lemma 3 establishes that there is no other Pareto-efficient allocation that intensity-dominates $x^{\prime}$ either. Thus, $x^{\prime}$ is intensity-efficient. ∎
@@
$\begin{array}[]{lllllllll}s^{\prime}_{1}(a^{\prime}_{i},a^{\prime}_{i+1})&>&s^{\prime}_{2}(a^{\prime}_{i},a^{\prime}_{i+1})&>&\cdots&>&s^{\prime}_{|N^{\prime}|}(a^{\prime}_{i},a^{\prime}_{i+1})&&\text{for all }i\leq|A^{\prime}|-1\end{array}$
Now recall that the allocation $x^{\prime}$ constructed above is intensity-efficient. This assigns $x^{\prime}_{m}=a^{\prime}_{m}\in A^{\prime}$ to each $N^{\prime}_{m}\in N^{\prime}$. Since $(N^{\prime},A^{\prime})$ is strictly monotonic but not homeo-monotonic, there are $N^{\prime}_{j},N^{\prime}_{l}\in N^{\prime}$ such that $N^{\prime}_{j}\trianglerighteq^{\prime}N^{\prime}_{l}$, jl>1j-l>1 and $s^{\prime}_{l}(a^{\prime}_{j},a^{\prime}_{l})>s^{\prime}_{j}(a^{\prime}_{j},a^{\prime}_{l})$. Since $(x^{\prime}_{j},x^{\prime}_{l})=(a^{\prime}_{j},a^{\prime}_{l})$, this implies that the allocation $x^{\prime\prime}$ which is identical to $x^{\prime}$ except that $(x^{\prime\prime}_{j},x^{\prime\prime}_{l})=(a^{\prime}_{l},a^{\prime}_{j})$ intensity-dominates $x^{\prime}$, contradicting the postulated intensity-efficiency of the latter allocation. $\square$
Proof of Theorem 4. To ease notation, we write $r_{l}$ instead of $r_{l}^{S}$ in much of the proof. We start with the following lemma, which is valid regardless of whether the intensity relation $\dot{\succsim}_{l}$ is linear or whether it induces a strict or weak preference relation on $A$:
Lemma 5.